A point charge q1 = +3.90 is at origin. How far should the second point charge of +6.30  be placed to have electric potential energy of 0.700 J?



1
Expert's answer
2022-03-30T13:51:05-0400

Potential energy

"U=\frac{kq_1q_2}{r}"

"r=\frac{kq_1q_2}{U}"


"r=\frac{9\times10^9\times3.90\times10^{-6}\times6.30\times10^{-6}}{0.700}=0.3159m"